### v1

# “It Explains What I am Currently Going Through Perfectly to a Tee”: Understanding User Perceptions on LLM-Enhanced Narrative Interventions

Abstract.
Stories about overcoming personal struggles can effectively illustrate the application of psychological theories in real life, yet they may fail to resonate with individuals’ experiences. In this work, we employ large language models (LLMs) to create tailored narratives that acknowledge and address unique challenging thoughts and situations faced by individuals. Our study, involving 346 young adults across two settings, demonstrates that LLM-enhanced stories were perceived to be better than human-written ones in conveying key takeaways, promoting reflection, and reducing belief in negative thoughts. These stories were not only seen as more relatable but also similarly authentic to human-written ones, highlighting the potential of LLMs in helping young adults manage their struggles. The findings of this work provide crucial design considerations for future narrative-based digital mental health interventions, such as the need to maintain relatability without veering into implausibility and refining the wording and tone of AI-enhanced content.
Story, Narrative, Digital Mental Health, LLM, Intervention
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXccs: Human-centered computing Empirical studies in HCI
1. Introduction
Stories about individuals overcoming personal struggles can serve as a powerful medium for promoting wellbeing (Bhattacharjee et al., 2022; Anderson and Wallace, 2015; Llewellyn-Beardsley et al., 2019; Kornfield et al., 2022b). By engaging with narratives that mirror their own experiences, individuals can draw parallels between their lives and those of the characters, fostering a space for self-reflection (Ferguson and Musheno, 2000). For instance, a narrative about someone overcoming disappointment from job search rejection could provide valuable insights to others facing similar setbacks. This reflection process can help readers recognize patterns in their behavior or thinking, leading to practical revelations that empower them to make positive changes (Bhattacharjee et al., 2023a, 2022). Moreover, engaging with such stories can validate and normalize the mental health challenges that individuals face (Kornfield et al., 2022a).
Given their substantial benefits, stories are increasingly being integrated into digital mental health (DMH) interventions (Meyerhoff et al., 2024; Bhattacharjee et al., 2022; Baumel et al., 2018). However, a significant challenge that many DMH interventions face, including those using storytelling, is their inability to deliver content tailored to the specific challenges that individual users encounter (Matthews and Rhodes-Maquaire, 2024). Many existing DMH interventions cannot be heavily tailored to users’ unique situations, instead relying on pre-defined rules to make coarse content adjustments that may not always resonate (Bhattacharjee et al., 2023b). This misalignment can cause the content to appear generic and disconnected from users’ immediate needs. Such perceptions not only increase the likelihood that the content will be overlooked but also risk dismissing users’ genuine struggles. Consequently, seemingly irrelevant narratives may undermine the potential effectiveness of a DMH intervention (Bhattacharjee et al., 2022).
Large language models (LLMs) are rapidly gaining recognition for their capabilities as personalized intervention tools and supportive scaffolds in a variety of settings (Bhattacharjee et al., 2024b; Kim et al., 2023; Abd-Alrazaq et al., 2023). These models excel in processing open-ended user inputs, enabling them to dynamically adapt and respond to a wide range of queries. Given their flexibility, LLMs could be tools that interpret users’ descriptions of their situations in order to create stories specifically tailored to one’s circumstances. We posit that such capability will significantly enhance the personalization of digital narrative interventions, making content more relevant and impactful. Revisiting the previous example of job search rejection, a user might gain more benefits from a narrative tailored to reflect the competitive nature of their field or environment. Alternatively, the story could address a specific cognitive distortion where they see setbacks in their job search as an ongoing pattern of defeat, failing to recognize their successes or factors out of their control. By connecting the story to specific circumstances and concerns, a narrative might provide deeper insights and more appropriate strategies, effectively addressing the user’s comprehensive experience.
Although LLMs offer these promises for generating personalized stories, implementing them effectively and reliably is non-trivial. They are not innately trained and lack the clinical and experiential knowledge to generate stories on their own (Bhattacharjee et al., 2024b). Considering the sensitivity of mental health struggles, applications of LLMs must avoid generating content that is harmful or insensitive to users’ needs (Demszky et al., 2023). To mitigate these risks, one approach could involve preloading LLMs with narratives and mental health content from previous DMH interventions, limiting them to generate content constrained within those boundaries. This would allow personalization to target changing the context of the story to match an individual’s circumstances while keeping the core takeaways and guidance in line with validated approaches.
In this work, we investigated how users perceive various aspects of LLM-enhanced stories designed to assist individuals in managing negative thoughts. To ensure that LLMs produce appropriate stories, our strategy integrated narratives gathered from real-life experiences that were curated by experts and previously utilized in DMH interventions (Meyerhoff et al., 2024). These stories were accompanied by prompt engineering emphasizing the intended qualities and structure of the stories to be generated. Our aim was to observe how individuals perceive these LLM-enhanced narratives and gather feedback on story qualities, such as the authenticity of the narratives and how effectively key takeaways were communicated. Additionally, we assessed the impact of these stories on changing participants’ beliefs in their negative thoughts and the intensity of their emotions.
Our research focused on young adults aged 18–25, a demographic particularly susceptible to mental health issues due to the inherent transitional nature of this life stage, encompassing significant changes in education, career, and personal relationships (Kornfield et al., 2022a). Statistically, this age group exhibits a much higher prevalence of mental health concerns compared to other adult groups (National Institute of Mental Health (2020), NIMH). These factors make them a relevant group for our research, where we aim to design stories that can help manage common challenges they face.
Our work was guided by the following research questions:
  • RQ1: How do young adults perceive the key narrative qualities of LLM-enhanced stories, such as authenticity and the ability to communicate takeaways, compared to human-written stories from past DMH interventions?
  • RQ2: How effective are LLM-enhanced stories as interventions for managing negative thoughts, such as reducing belief in negative thoughts and inspiring solutions, compared to human-written stories from past DMH interventions?
To address these research questions, we conducted a randomized experiment with two groups of young adult participants: crowdworkers ($N=174$) and students ($N=172$). All participants engaged in a survey activity where they were randomly assigned to interact with either a human-written story from past interventions (Non-LLM story) or an LLM-enhanced story. Our findings revealed that LLM-enhanced stories were perceived as comparably authentic to Non-LLM stories, yet they performed better on average in communicating key takeaways, promoting reflection, reducing belief in negative thoughts, and inspiring actionable solutions. Our results suggest the promise of LLMs to create custom narratives supporting personal wellbeing, and inspire design considerations for LLM-enhanced narrative interventions, such as the need for a careful balance between relatability and avoiding implausibility and the importance of refining the tone and wording in AI-generated content.
2. Related Work
For the review of related work, we begin by discussing how story-based interventions can promote psychological wellbeing. Following this, we talk about the need for dynamic content generation in DMH interventions, focusing on the challenges and potential of LLMs in this context.
2.1. Promoting Psychological Wellbeing Through Stories
Storytelling has long been a traditional method for imparting morals and lessons, as evidenced through various mediums such as fables, allegories, parables and poems (Sarlej and Ryan, 2012; Gibbs et al., 2002). Among the most famous examples of moral storytelling are Aesop’s Fables (Aesop, 1998), famous for concluding with succinct moral takeaways like “look before you leap” or “honesty is the best policy.” Recent research underscores the role of storytelling in supporting psychological wellbeing, highlighting how narratives can make challenging experiences more relatable and offer paths to recovery (Nurser et al., 2018; Llewellyn-Beardsley et al., 2020, 2019; Anderson and Wallace, 2015). Through these stories, individuals can find inspiration to mirror the story characters’ resilience and recovery strategies in their own lives (Shaw and Homewood, 2015; Nurser et al., 2018; Williams et al., 2018; Lipsey et al., 2020). Furthermore, storytelling can encourage self-reflection, helping individuals see parallels between themselves and the characters, which can lead to a deeper understanding of their mental state and encourage constructive changes in thought patterns and behaviors (Dym et al., 2019; Payne, 2006; Hinyard and Kreuter, 2007; Bhattacharjee et al., 2022). This reflective practice is also crucial in various therapeutic settings where sharing and discussing personal stories about one’s journey through adversity helps individuals learn from each other (Payne, 2006; Lipsey et al., 2020).
Storytelling serves as one method of fostering “diffuse sociality,” (Burgess et al., 2019) a low-effort way for individuals to connect with others and share experiences without direct or face-to-face interaction. This form of connection can significantly reduce feelings of isolation by revealing that others face similar struggles (Burgess et al., 2019; O’Leary et al., 2017). Such a sense of connectedness can be especially beneficial in DMH interventions, where users may prefer to engage independently. Research shows that individuals can feel a strong connection to the characters of a story, seeing their own challenges reflected in the narratives, and finding reassurance in shared human experiences (Bhattacharjee et al., 2022; Akram et al., 2020; Lipsey et al., 2020).
There is growing evidence supporting the benefits of narrative approaches in promoting psychological wellbeing and behavioral change (Bhattacharjee et al., 2022; Saksono et al., 2021; Grimes et al., 2008). Research has shown that storytelling interventions can reduce symptoms of depression across various populations (Vromans and Schweitzer, 2011; Rodríguez Vega et al., 2014); for instance, Rodríguez Vega et al. (2014) have utilized mindfulness-based narrative therapy as a therapeutic intervention for treating depression in cancer patients. Behavioral change interventions have leveraged stories to subtly guide audiences toward specific viewpoints or actions by immersing them in relatable narratives (Bilandzic and Busselle, 2013). Some interventions embed educational content within an entertainment framework containing narratives, engaging audiences while delivering important messages (Singhal and Rogers, 2012). Stories have been used successfully to encourage physical activity (Saksono et al., 2021), promote healthy eating habits (Grimes et al., 2008), and counter stigma against mental health conditions (Hospital, 2022; Evans-Lacko et al., 2014; Shoes, 2022) and sexual harassment victims (Dimond et al., 2013). This broad applicability underscores the power of stories to inspire behavioral change and foster psychological wellbeing in diverse contexts.
2.2. Dynamic Content Generation in DMH Interventions
Because mental health challenges are deeply personal and varied, static or pre-generated DMH content often fails to resonate with users (Bhattacharjee et al., 2024a). Dynamic content generation addresses this shortcoming by producing personalized and contextually relevant outputs (Datta et al., 2004). In the context of DMH interventions, this capability is critical for enhancing engagement, effectiveness, and user satisfaction (Bhattacharjee et al., 2023b; Baumel et al., 2019). However, delivering these benefits remains challenging.
2.2.1. Challenges of Dynamic Content Generation:
Early DMH interventions, such as rule-based chatbots, relied on pre-scripted responses and decision trees for support (Weizenbaum, 1966; Abd-Alrazaq et al., 2019). While these systems have been effective at generating structured conversations, they have been inherently limited in their capacity for dynamic content generation. Since responses were predefined, the systems struggled to adapt to the unique and evolving contexts of individual users’ life experiences (Abd-Alrazaq et al., 2019; Bhattacharjee et al., 2023b). As a result, the interactions often felt static and generic, lacking the personalization necessary to address users’ specific needs in real time.
Some prior studies (Ma et al., 2023; Morris and Picard, 2014; Smith et al., 2021; Saksono et al., 2023; Morris et al., 2015) explored supporting user wellbeing through dynamic content generated via crowdsourcing. For instance, Panoply (Morris and Picard, 2014; Morris et al., 2015), a crowd-powered system, allowed users to submit descriptions of their stressors. These descriptions were then sent to crowdworkers who crafted supportive messages in response. Similarly, Flip*Doubt was a crowd-powered web application that provided users with cognitive reappraisals of negative thoughts (Smith et al., 2021). Additionally, many online peer support groups have been established with a similar motivation (Baumel et al., 2018; Wang et al., 2023). These platforms aim to foster wellbeing by connecting individuals facing similar challenges, enabling them to share stories, learn from each other, and engage in discussions that promote reflection. However, these platforms heavily depend on the availability and responsiveness of participants, which is not always guaranteed. Moreover, the varying levels of comfort among crowdworkers and peers in sharing their experiences could limit the effectiveness of the support provided. Furthermore, crowdworkers and peers may also require training to ensure they deliver appropriate and empathetic support (Chen et al., 2021).
The advent of machine learning and deep learning brought some advancements in dynamic content generation for DMH interventions. These models began to harness personalized data, such as user history, mood patterns, and behavioral cues, to generate responses that were more attuned to each user (Torous et al., 2018; Adler et al., 2021; Fitzpatrick et al., 2017). Numerous conversational chatbots were developed during this era, using AI algorithms to choose from pre-set branches (Fitzpatrick et al., 2017; Meyerhoff et al., 2024; Aguilera et al., 2020), but their capacity to process open-ended input from users and adapt accordingly was still very limited. Natural language processing (NLP) techniques, like information retrieval or word embedding (Carpenter et al., 2016), played important role in improving how systems interpreted free-form user input, enabling more relevant and context-sensitive outputs (Le Glaz et al., 2021). For instance, Morris et al. (2018) have employed these techniques to repurpose responses from existing online peer support databases, presenting the content that most align with their specific concerns. However, despite these advancements, the ability to dynamically generate new content still remained limited. These systems excelled at pattern recognition but fell short in the creativity needed to produce truly customized content without large-scale training datasets. As a result, most systems continued to rely on selecting responses from a predefined content bank rather than creating entirely new content (Abd-Alrazaq et al., 2019).
2.2.2. Potential of LLMs in Dynamic Content Generation for DMH Interventions
Recent advancements in the development of LLMs like GPT and Gemini represent a leap forward for dynamic content generation. These models have the potential to generate coherent and rich text, offering a significant improvement over earlier systems that relied on selecting from predefined content (Bhattacharjee et al., 2024b). These capabilities can extend to various domains, particularly in DMH interventions, where understanding nuanced user needs is crucial. LLMs have the potential to tailor interventions to an individual’s unique experience that might involve a challenge they face or a recent experience. This is particularly facilitated by the LLMs’ ability to process open-ended input, allowing users to express and respond to their situations in rich detail and in any form they choose.
For example, Sharma et al. (2023b) developed an LLM-based cognitive reframing tool that dynamically adapts its responses based on the specific challenges faced by users. Similarly, CLOVA CareCall, a conversational AI, collects data on individuals’ general health and engages as a conversational partner to alleviate loneliness by generating dynamic, human-like questions and responses (Jo et al., 2023, 2024). Additionally, Hedderich et al. (2024) introduced an LLM-based chatbot designed to assist teachers in providing support to adolescents dealing with cyberbullying, with responses that recognize and adapt to the context of each situation. Studies such as the one by Kim et al. (2024) have utilized LLMs to offer contextual writing support in journal writing. In their work, Kim et al.’s system provided real-time assistance as users expanded their journal entries. Collectively, these applications underscore the promise of LLMs in generating dynamic support that not only acknowledges but also responds appropriately to the nuanced needs of users.
However, adapting LLMs for dynamic content generation in DMH tools is not a straightforward task. Research has shown that LLMs can produce inaccurate or clinically invalid suggestions within the wellbeing context (Ma et al., 2024; Demszky et al., 2023). This highlights the need for design considerations to ensure that content is not only accurate but also delivered effectively to support wellbeing. Additionally, it is crucial to determine which intervention elements can be personalized and which must remain standardized. Bhattacharjee et al. (2024b) emphasize that LLMs should operate within specific evidence-based psychological strategies when providing support in managing emotions. In this domain, LLMs require careful direction regarding the situational context of the user and should adhere strictly to predetermined structures and content guidelines.
3. Study Design
Below, we describe the design of our study and the various steps we employed to execute it.
3.1. Procedure
Participants were recruited for a brief activity, lasting between 10 and 20 minutes, aimed at aiding the development of technology-based narrative interventions for managing negative emotions. During recruitment, it was communicated that their participation would contribute to research in this area. The survey activity, which included dynamic story selection and generation through the use of LLMs, was conducted on the Qualtrics platform.
After providing affirmative informed consent, participants were asked to select a common challenging thought that they frequently encounter (we label this question as Q-Challenging-Thought). This selection was made from six options identified from previous research on mental health challenges in young adults (Meyerhoff et al., 2024). The options included feelings of being down on oneself, worrying about the future, feelings of struggling, feelings of underachievement compared to others, disappointments in reliance on others, and social anxiety. Participants then evaluated their belief in, and the emotional intensity of, the selected thought on a scale from 1 (’very low’) to 7 (’very high’), inspired by questions from a past study (Sharma et al., 2023b). They were also requested to provide a detailed description of their thought, circumstances, and emotions to the extent they felt comfortable (we label this question as Q-Challenging-Thought-Description).
Following this initial sequence of questions, participants were randomly assigned to engage with either a story sourced from prior studies (Non-LLM condition) or an LLM-enhanced version of the story (LLM condition). In both groups, stories addressed the same challenging thought endorsed by the participant. Participants were blind to the condition they were assigned to; in both conditions, stories were described broadly as depicting the experience of someone else going through a similar situation. Each story concluded with a reflection question, prompting participants to reflect on their behavior and potential strategies to overcome their challenges. The number of words written by participants in this reflection was captured as an indicator of their engagement with the intervention (Wang et al., 2023). We detail the design of stories in both conditions in Section 3.3.
Subsequently, participants rated four qualities of the stories — such as authenticity and the effectiveness of communicating key takeaways — on a scale from 1 (’very low’) to 7 (’very high’). They also provided qualitative feedback on their ratings and additional quantitative evaluations concerning related qualities like perceived relatability and engagement. Qualitative comments on both helpful and unhelpful aspects of the stories were also collected. These questions were also inspired from prior literature (Bhattacharjee et al., 2022, 2024a; Sharma et al., 2023b). Following best practices in survey design, we also incorporated attention-check questions in the middle of the survey to ensure participant attentiveness and the quality of the responses (Hauser and Schwarz, 2016).
Table 1 highlights the quantitative questions, and we label them for future reference. All questions were asked on a 1–7 scale, where 1 represents the lowest or least favorable response (e.g., “strongly disagree” or “not at all”), and 7 represents the highest or most favorable response (e.g., “strongly agree” or “extremely”). We also highlight the sources of these questions, along with the corresponding research question.
Table 1. Quantitative questions from the survey
Questions asked before and after reading the story
Question Label Actual Question Corresponding RQ Source
Belief in Negative Thought How strongly do you believe in your thought? RQ2 (Sharma et al., 2023b)
Emotion Intensity How strong is the emotion associated with this thought? RQ2 (Sharma et al., 2023b)
Questions asked after reading the story
Question Label Actual Question Corresponding RQ Source
Perceived Authenticity How would you rate the authenticity of the story in portraying characters, settings, and emotions? RQ1 (Bhattacharjee et al., 2022)
Perceived Ability to Communicate Key Takeaways How effectively do you think the story communicated its key messages or lessons? RQ1 (Bhattacharjee et al., 2022)
Perceived Ability to Promote Reflection How much did the story encourage you to reflect on your own thoughts, experiences, or beliefs? RQ1 (Bhattacharjee et al., 2022)
Perceived Balance of Positivity with Realistic Struggle How well did the story balance positive outcomes with realistic portrayals of challenges? RQ1 (Bhattacharjee et al., 2022)
Perceived Relatability Agreement with the statement: ‘I could relate to the experiences shared in the story.’ RQ2 (Meyerhoff et al., 2024)
Perceived Capacity of Inspiring Solutions Agreement with the statement: ‘The story helped me think of potential ways I can manage negative thoughts.’ RQ2 (Bhattacharjee et al., 2024a)
Likelihood of Future Engagement Agreement with the statement: ‘I would love to read similar stories in future.’ RQ2 (Bhattacharjee et al., 2024a)
For our quantitative comparison, we hypothesized that participants would perceive the LLM-enhanced stories more favorably across all qualities except for authenticity, compared to those in the Non-LLM condition. This hypothesis was based on the LLMs’ potential to improve DMH content across various settings. However, we expected lower scores for authenticity in the LLM condition because these stories, while based on human-written content, were adapted by LLMs, unlike the stories in the Non-LLM condition.
3.2. Participants
We conducted two separate experiments to compare stories in the LLM and the Non-LLM conditions, recruiting crowdworkers from the Prolific platform and students from a large computer science course at a major North American University. All participants were required to be between 18 and 25 years old. For the crowdworkers, an additional requirement was residency in North America. Since our study aimed to help the general population manage common challenging thoughts, we did not impose any criteria related to wellbeing or mental health symptoms.
A total of 180 crowdworkers and 179 students initially participated in our study. However, data from 6 crowdworkers and 7 students were discarded due to quality issues identified through instructional manipulation checks of attention-check questions  (Oppenheimer et al., 2009), resulting in a final sample of 174 crowdworkers and 172 students.
Participants were randomly assigned to either the LLM or the Non-LLM condition. After the assignments, we had 87 crowdworkers in both the Non-LLM and LLM conditions, and 90 students in the Non-LLM condition and 82 in the LLM condition. We refer to them using a combination of the first letters of their group (Crowdworker = C, Student = S) and their condition (Non-LLM = N, LLM = L), followed by a numeric ID indicating their participant number within that condition. For example, CL82 refers to a crowdworker in the LLM condition, while SN56 refers to a student in the Non-LLM condition. Crowdworkers were compensated at a rate of \$10 USD per hour for their participation, while student participants received a 1% credit towards their course grade. Table 2 details participants’ demographic information, including age, gender, and race.
Table 2. Demographics of crowdworkers and students
Crowdworkers (N = 174) Students (N = 172)
Age, mean ± std err 22.5 ± 0.2 20.9 ± 0.1
Gender, N (%)
Women 101 (58.0%) 25 (14.5%)
Men 63 (36.2%) 142 (82.6%)
Non-binary/third gender 7 (4.0%) 0 (0%)
Self-described 1 (0.6%) 2 (1.2%)
Not reported 2 (1.1%) 3 (1.7%)
Race, N (%)
White 68 (39.1%) 19 (11.0%)
More than one race 13 (7.5%) 6 (3.5%)
Black or African American 13 (7.5%) 6 (3.5%)
Asian 65 (37.4%) 126 (73.3%)
American Indian or Alaskan Native 2 (1.1%) 0 (0%)
Not reported 13 (7.5%) 15 (8.7%)
3.3. Design of the Stories
In this section, we explain the design of stories in the Non-LLM and the LLM conditions. All of the stories in the Non-LLM condition, accompanied by an example LLM-enhanced version, are provided in Appendix A.
3.3.1. Design of Stories in the Non-LLM Condition
In the Non-LLM condition, our stories are sourced from Meyerhoff et al. (2024), who themselves based their stories on the experiences reported by young adult participants according to Kornfield et al. (2022a). To generate stories, Crowdworkers were given a list of common challenges to write about, or they could supply their own. They were then asked how they took a step that helped them better manage the challenge or feel better about it, even if it was a small or incremental step. They were provided with examples of evidence-based psychological strategies they may have used in addressing their challenge, or they were allowed to describe their own approach. Crowdworkers were asked to write true stories in the first person and to include as much detail as they could about their challenge and how they had addressed it, all while omitting identifying information. After clinical psychology and human-computer interaction researchers edited and curated the stories for clarity and to align with the intended psychological strategies, Meyerhoff et al. (2024) eventually deployed these stories as a part of a text messaging program designed to help young adults manage mental health and wellbeing.
Each story corresponded to one of the six challenging thoughts outlined in Q-Challenging-Thought, illustrating how a young adult navigates and overcomes such challenges. These stories were designed to promote evidence-based psychological strategies for managing negative thoughts, such as practicing self-compassion (Allen and Leary, 2010) or engaging in cognitive restructuring (Clark, 2013). The stories featured four essential qualities as outlined by Bhattacharjee et al. (2022): authenticity, key takeaways, active participation (i.e., promoting reflection), and a balance between positivity and realistic struggles. We describe below how these stories were crafted to embody each desired quality:
  • Authenticity: Crafted from a first-person perspective, the stories aimed to endow the narrative character with a distinct voice, facilitating a deeper connection between the character and the reader (Delgado, 1990). By incorporating detailed descriptions of the characters and their circumstances, the stories also created vivid and relatable imagery for the audience (Rosenthal, 1971). For instance, instead of generalizing about high school graduates, they mentioned specifics such as a high school graduate feeling overshadowed by peers who have seemingly achieved more or dealing with intense self-criticism while working from home during the pandemic.
  • Key Takeaways: Drawing from the format of Aesop’s Fables (Aesop, 1998), each story concluded with a clear message for the reader, linking a psychological principle to the narrative’s characters and events. For example, stories included takeaways like “Being deliberate about the things we tell ourselves can make a difference in how we feel.” or “Self-compassion is all about giving the same kindness to ourselves we give to others.”.
  • Promoting Reflection: The stories fostered interactive engagement by including a reflective question at the end of each story, such as “Have you ever faced a similar challenge?”. This technique was designed to prompt users to evaluate their own experiences in light of the story’s message, facilitating personal reflection. This method is akin to the dialogic inquiry used in storytelling, where the storyteller periodically pauses to invite listeners to relate the narrative to their own lives (Wells, 2000; Slovák et al., 2016).
  • Balance Between Positivity and Realistic Struggles: Each story ended on a positive note, depicting an individual making progress or gaining a new perspective on their challenges and illustrating the value of change. The conclusions were deliberately realistic; rather than depicting characters who have made enormous progress or solved their problems entirely, the stories emphasized that small, incremental steps can lead to meaningful improvements in one’s situation (Bhattacharjee et al., 2022).
3.3.2. Design of Stories in the LLM Condition
Our implementation of LLM-enhanced narratives was informed by recent research across various contexts, which demonstrated the benefits of adapting content to align with user contexts. (Bhattacharjee et al., 2024b; Stade et al., 2024). For our project, we employed a prompt engineering method. The prompt design was an iterative effort involving faculty and graduate students with expertise in human-computer interaction, LLMs, psychology, cognitive science, and DMH. The prompt design underwent multiple rounds of revision, applying the prompts to a variety of stories from past interventions and refining them iteratively.
Following recommended practices of prompt engineering (Stade et al., 2024; Ekin, 2023), we developed a structured prompt using OpenAI’s GPT-4, incorporating example stories from the Non-LLM condition to ensure adherence to the structured format discussed in Section 3.3.1. The prompt was designed to weave in three responses from participants:
  • Selected option from Q-Challenging-Thought
  • Response to the open-ended question Q-Challenging-Thought-Description
  • Corresponding story to the selected option from Q-Selected-Thought
We also included instructions to avoid the introduction of new suggestions or strategies not present in the examples. The stories were designed to maintain a similar length to those in the Non-LLM condition ($\leq$250 words) and concluded with a reflective question. The complete prompt is as follows:
{mdframed}
[backgroundcolor=gray!20]Given the user’s thought, [[selected choice from Q-Challenging-Thought]], and their provided description, [[response to the open-ended question Q-Challenging-Thought-Description]], your task is to share a narrative. This narrative should recount the journey of an individual who encountered challenges similar to yours and the steps they took towards making progress, incorporating relevant examples and strategies for improvement. Here are some guidelines to follow:
Authenticity: The experience shared should feel genuine and relatable, mirroring real-life emotions and scenarios. The story should be told from a first-person perspective.
Key Takeaways: Clearly articulate lessons or insights that the user can apply to similar situations in their life.
Reflective Question: At the end of the narrative, there should be a reflective question along the lines of “Have you had a situation where you had to deal with a similar challenge?” or ”What strategies might work during such challenging times?”.
Balancing Positivity with Realistic Struggles: While maintaining a positive outlook, acknowledge the challenges and struggles inherent to the situation.
Use the following example as a basis for your narrative. It is important that you do not introduce new suggestions or strategies beyond those mentioned in the example.
[[Corresponding story to the selected choice from Q-Selected-Thought]]
Please generate a first-person narrative that incorporates these guidelines and examples, making it as engaging and helpful as possible for the user to gain new perspectives and coping mechanisms. The text should start with “Here’s someone’s experience that might resonate with you: [[newline]]” and should not contain more than 250 words. As shown in the example, conclude with a clear separator, “******,” followed by a prompt for a reflective question that starts with “Please answer this reflective question:”
3.4. Data Analysis
We determined the statistical significance of differences between the groups for each measure using one-tailed Welch’s t-tests. P-values were adjusted using the Benjamini-Hochberg (BH) procedure for multiple comparisons. The null hypothesis (H0) posited that the mean for the hypothesized superior group is less than or equal to the mean for the other condition. Conversely, the alternative hypothesis (Ha) contended that the mean for the hypothesized superior group is greater than that of the baseline condition.
After the data were anonymized and cleaned, we utilized the thematic analysis approach outlined by Clarke and Braun (2017) to examine the qualitative data. First, two coders from the research team reviewed the data to familiarize themselves with the data. Using an open-coding method (Khandkar, 2009), each coder independently developed an initial codebook for the data of 25 participants. We then implemented a consensus coding method similar to that outlined by Clarke and Braun (2017). During this phase, the coders held several joint sessions to review and integrate their separate codebooks. This consensus coding process, through repeated discussions and refinement, was designed to improve reliability and minimize personal biases. The coders determined common codes, refined the definitions, and removed codes that did not directly relate to the research questions. The updated codebook was then applied iteratively to the data from an additional 25 participants, making adjustments as new findings emerged. After finalizing the codebook, each coder independently applied the established codes to half of the remaining data.
Our analysis generated 138 codes, capturing themes such as “new perspective,” “matching personal struggles,” “robotic,” and “clear takeaways.” Coders used axial coding (Williams and Moser, 2019) in subsequent meetings to organize these codes into major themes. Detailed findings structured by these themes are presented in Section 4.
3.5. Ethical Considerations
Our research activity was approved by the Research Ethics Board at the first author’s university. Given the nature of our research, which involved aspects of reflection on negative thoughts and challenging life situations, we were mindful of the ethical considerations at play throughout the study, and we took proactive measures to address potential concerns. All participants provided affirmative informed consent prior to beginning any study procedures. They were informed about the possibility that participation might cause anxiety, sadness, or other forms of distress. Participants were also requested to not include any identifiable information in their responses to ensure confidentiality and privacy.
Moreover, while we did not ask participants about their experiences with suicide-related thoughts and behaviors, we recognized that suicidal thoughts, behaviors, or other risk-related issues may arise due to the open-ended nature of the questions and assessments as well as the probabalistic nature of LLM responses. We provided all participants with contact information for crisis services, such as suicide prevention helplines, to ensure they had access to support, if needed. Additionally, the research team reviewed all survey responses daily during the study period to monitor for any indications of suicide-related risk, such as suicidal ideation or self-harm. Team members were trained to respond appropriately, and assess participants using the Columbia-Suicide Risk Assessment protocol (Posner et al., 2008) if any potential suicide or self-harm-related responses were identified. However, no such risks emerged during the course of the study, and we did not need to conduct any follow-up assessments.
4. Results
We structure our findings around the research questions we originally set. First, we examine how LLM-enhanced stories were perceived in terms of exhibiting desirable narrative qualities. Subsequently, we explore their effectiveness as interventions, specifically looking at their impact on managing negative thoughts. We primarily explore user perceptions of stories in the LLM condition, occasionally drawing comparisons to the Non-LLM condition to highlight differences and similarities.
4.1. LLM-Enhanced Stories in Exhibiting Key Narrative Qualities
Table 3 presents the summary statistics of the ratings given by participants regarding the narrative qualities of the stories. We observed that the stories from the LLM condition were perceived to be superior in all the narrative qualities other than authenticity. The differences between groups in authenticity ratings were not significant. The remaining differences were statistically significant among the crowdworkers, whereas among the students, only the difference in the perceived ability to promote reflection was significant. Below, we support these trends using insights drawn from participants’ comments.
Table 3. Summary statistics and measures of the stories’ narrative qualities with and without LLM enhancement. The results are separated across the two populations that were studied: crowdworkers and university students.
Crowdworkers ($\textbf{n}_{Non-LLM}=\textbf{n}_{LLM}=87$) Students ($\textbf{n}_{Non-LLM}=90,\textbf{n}_{LLM}=82$)
Mean Mean
Quality Hypothesized Superior Group Non-LLM LLM p-value Non-LLM LLM p-value
Perceived Authenticity Non-LLM 5.03 $\pm$ 0.17 4.92 $\pm$ 0.17 0.31 4.34 $\pm$ 0.18 4.29 $\pm$ 0.19 0.42
Perceived Ability to Communicate Key Takeaways LLM 5.51 $\pm$ 0.14 6.01 $\pm$ 0.10 ¡0.01** 5.27 $\pm$ 0.16 5.48 $\pm$ 0.17 0.21
Perceived Ability to Promote Reflection 4.71 $\pm$ 0.20 5.31 $\pm$ 0.17 0.02* 3.96 $\pm$ 0.20 4.55 $\pm$ 0.20 0.04*
Perceived Balance of Positivity with Realistic Struggle 4.80 $\pm$ 0.16 5.38 $\pm$ 0.16 ¡0.01** 4.29 $\pm$ 0.18 4.62 $\pm$ 0.17 0.15
* p ¡.05, ** p ¡ .01, *** p ¡ .001
4.1.1. Perceived Authenticity
Participants in both LLM and Non-LLM conditions were drawn to the authenticity of their stories because of their relatability and the sincerity depicted in the characters’ journeys. These narratives were appreciated for their candid portrayal of often unspoken feelings, aligning closely with personal and broadly shared human experiences.
However, for some participants in the Non-LLM condition, the very universality that made the stories relatable also rendered them suspect in terms of authenticity. They perceived these universal themes as too generic, suggesting a lack of specificity that could make the stories feel impersonal.
Participants in the LLM condition provided various insights into the factors contributing to perceived authenticity in the LLM-enhanced stories. SL473 remarked on the relatability and credibility of their story, stating:
I think the setting and thoughts portrayed in the story felt very natural and authentic as it represents a common feeling and scenario that many people face. The fact that the “journey still wasn’t smooth all the time” helped to make the story feel more realistic.
Other participants resonated with their stories as reflective of their own potential future experiences or as echoing a shared journey with others facing similar challenges. The familiarity of the scenarios and the realistic solutions proposed were key elements that the participants found compelling.
The quality of the writing also played a crucial role in enhancing the authenticity of the stories. Descriptive and vivid prose helped participants clearly visualize the scenarios, fostering a deeper connection with the content. CL73 shared, “I could picture the scenario pretty well in my mind and resonate with it.” However, several participants criticized the tone and writing style of their LLM-enhanced stories as being too mechanical and impersonal. Doubts about AI involvement were raised due to perceived shortcomings in the depth of the protagonists and their emotions. In fact, some participants commented that their stories shared patterns similar to the typical responses made by ChatGPT. CL41 noted:
Anyone could have written it, and the sentence structure is very even and AI-like. I have used a lot of AI text generators in the past so it gives me the same feeling. …[[The story was]] generic and removed from their own feelings.
Furthermore, some participants felt that the LLM-enhanced stories were overly tailored to their specific situations, making the AI’s involvement too apparent and diminishing the authenticity of the narratives. This precise mirroring of their inputs led to a perception that the stories were merely processing and regurgitating information, which reduced their impact and made them feel less genuine.
4.1.2. Perceived Ability to Communicate Key Takeaways
Participants in both conditions valued the structured progression within their stories, which clearly delineated the flow from beginning to end. However, some participants in the Non-LLM condition commented that the takeaways were not very evident as they did not relate to their stories. For instance, SN63 acknowledged understanding the message but felt a disconnect from their personal experiences.
Participants in the LLM condition felt that their stories were crafted to closely mirror their current life situations, which made the messages more poignant and relevant. This led CL34 to say, ”It explains what I am currently going through perfectly to a tee.” Customization helped participants see their stories not just as general advice, but as directly applicable guidance. The stories’ structure played a crucial role in effectively communicating key takeaways as well. By charting a clear progression from past to future, the stories in the LLM condition provided contextual grounding that guided the reader through the speaker’s process of discovering solutions. CL19 explained:
It broke down the story clearly. Explained why this person felt this way and what circumstances were around it. Clarified the feelings they were having and the choices they made to better themselves and finished with how this struggle makes them a better person now.
Criticisms of the LLM-enhanced stories in this regard were minimal, mainly focusing on the tone and presentation of the content. The feedback echoed previous sentiments, highlighting that mechanical and impersonal tones detracted from the authenticity of the narratives. Additionally, there was some concern from the students that advice that seemed to be generated by AI might not be taken as seriously by readers, potentially undermining the impact and credibility of the guidance provided.
4.1.3. Perceived Ability to Promote Reflection
Participants in the Non-LLM condition valued how their stories depicted common experiences that reinforced actions and introduced fresh perspectives, but they were sometimes criticized for being too generic. This generality made their stories seem less reliable as catalysts for deep reflection. The lack of specific, detailed scenarios in these stories sometimes led participants to perceive the stories as insufficiently engaging or thought-provoking.
Participants in the LLM condition were more enthusiastic about the successes that their stories had in promoting reflection. They particularly noted the influence of the recommended behaviors within the stories. This perceived applicability allowed participants to see the advice and strategies presented as both relevant and practical, directly connecting their stories to their own life situations.
Reflection was triggered in various ways. For participants like CL1, the stories validated the idea that “everyone has different paths and progressions in their lives”, helping them appreciate and accept the diversity in personal journeys and perspectives. Additionally, these stories reinforced beliefs or approaches that participants were already exploring or implementing to manage negative thoughts. For some others, the situations depicted were relatable, but the recommended behaviors provided fresh perspectives and potential methods for participants to consider. CL38 expressed:
To be fair, the story did give me a step back to look at my situation from a different perspective. It made me think ‘Maybe I am a little too harsh on myself.’
While relatability was often beneficial, it sometimes impeded participants’ ability to reflect. In certain cases, participants had encountered similar scenarios repeatedly in the past, so the familiarity stripped away the novelty that might have prompted more profound contemplation about their own circumstances. CL65 mentioned that a slightly different situation in the story might have offered a fresh perspective on the problem. These findings highlight how overly familiar situations can diminish the potential for deeper reflection.
4.1.4. Perceived Balance of Positivity with Realistic Struggles
The structured progression of events in the stories was viewed favorably in both conditions, as it helped maintain a balance between positivity and realistic portrayals of struggles, anchoring the narratives and ensuring they resonated with the audience to some extent. However, participants in the LLM condition were particularly appreciative of the balance achieved between positivity and realistic struggles They particularly highlighted their stories’ empathetic acknowledgment of the protagonist’s struggles, which mirrored the challenges faced by the readers themselves. This recognition of shared difficulties established a foundation of empathy, making participants feel understood and more connected to the narrative. Additionally, their stories were commended for illustrating the small, incremental steps protagonists took to address and improve their challenges, enhancing their realism and applicability.
Participants appreciated that their stories did not promise unattainable perfection but instead portrayed a more truthful journey towards improvement. This realistic depiction resonated well with participants like SL86, who stated:
I felt like it was an incredibly realistic portrayal. At the end of the story, the storyteller even mentions that not everything went perfect, but they were able to face challenges more effectively. That outcome is realistic and something I hope to obtain in the future.
However, participants noted that both the Non-LLM and LLM-enhanced stories often adopted a more positive tone than they would have preferred. This sentiment was especially pronounced when participants perceived their situations as beyond their control, leading to a disconnect with the occasional overly optimistic narrative. CL35 expressed this concern:
It’s not that easy to just focus on what you have in control and the rest will follow in reality. When you have a lot of things on your plate it gets stressful to manage.
Participants suggested that in some difficult situations, particularly where people may feel they lack control over their situation, portraying progress might be seen as an oversimplification. They recommended acknowledging the challenges more explicitly in such situations and potentially ending the stories in a more open-ended manner, rather than always concluding with progress.
4.2. LLM-Enhanced Stories as an Intervention to Manage Negative Thoughts
Table 4 presents the summary statistics of the story attribute ratings given by participants. We observed that the stories from the LLM condition were perceived to be superior in all the qualities. Save for the ratings of perceived emotional intensity, the differences in ratings were statistically significant among the crowdworkers. Among the students, only the differences in perceived relatability, perceived capacity of inspiring solutions, and length of written reflection were statistically significant. Below, we supplement these findings with participant comments that highlight both the beneficial aspects and the areas for improvement of the stories.
Table 4. Summary statistics and measures of the stories’ intervention qualities with and without LLM enhancement. The results are separated across the two populations that were studied: crowdworkers and university students.
Crowdworkers ($\textbf{n}_{Non-LLM}=\textbf{n}_{LLM}=87$) Students ($\textbf{n}_{Non-LLM}=90,\textbf{n}_{LLM}=82$)
Mean Mean
Quality Hypothesized Superior Group Non-LLM LLM p-value Non-LLM LLM p-value
Perceived Reduction in Belief in Negative Thought LLM 0.26 $\pm$ 0.10 0.64 $\pm$ 0.13 0.02* 0.49 $\pm$ 0.12 0.65 $\pm$ 0.12 0.21
Perceived Reduction in Emotion Intensity 0.59 $\pm$ 0.13 0.89 $\pm$ 0.12 0.06 0.44 $\pm$ 0.10 0.65 $\pm$ 0.12 0.15
Perceived Relatability 5.16 $\pm$ 0.17 5.75 $\pm$ 0.13 ¡0.01** 4.48 $\pm$ 0.17 5.05 $\pm$ 0.18 0.03*
Perceived Capacity of Inspiring Solutions 4.74 $\pm$ 0.18 5.46 $\pm$ 0.16 ¡0.01** 4.14 $\pm$ 0.19 4.76 $\pm$ 0.18 0.04*
Likelihood of Future Engagement 4.53 $\pm$ 0.20 5.21 $\pm$ 0.17 ¡0.01** 3.92 $\pm$ 0.20 4.30 $\pm$ 0.20 0.15
Number of Words Written for Reflection 31.15 $\pm$ 2.28 44.22 $\pm$ 4.04 ¡0.01** 25.29 $\pm$ 2.42 33.56 $\pm$ 2.65 0.04*
* p ¡.05, ** p ¡ .01, *** p ¡ .001
4.2.1. Reevaluating Existing Thoughts and Beliefs
Participants in the Non-LLM condition found that their stories allowed them to organize and reevaluate their thoughts and emotions in a productive manner. Participants recognized the benefits of the reflective writing activity associated with these stories, demonstrating the potential of narratives in facilitating the process of reflection among participants.
Participants in the LLM condition expressed similar sentiments, noting that they were able to reevaluate their negative thought patterns from a more neutral point of view. For instance, CL11 said the following about reading a story related to a person still being in school while their friends had already completed their studies:
Now I understand that everybody has their own timing and even if it takes me more time to find out what I want, at least I’ll be happy and sure of what I want to do.
The reevaluation process also involved a reflective writing activity associated with the reflection question. As shown in Table 4, both crowdworkers and students wrote more words on average in response to the LLM-enhanced stories. This activity aimed to provide participants with an opportunity to organize their thoughts and emotions into written form, which may have helped them restructure their thoughts to be more helpful. SL28 expressed one such change, saying:
Despite all the negative thoughts I had about myself, I should be proud of the things that I have already achieved which was by completing university and getting that bachelor’s degree. …everyone should be enjoying the journey and not just the destination.
These findings suggest that both Non-LLM and LLM-enhanced stories could support participants in reevaluating their thoughts and beliefs, offering them new insights and perspectives.
4.2.2. Potential of Stories in Inspiring Solutions and Actions
Participants in the Non-LLM condition generally observed a lack of pragmatic solutions and actions in their stories compared to those in the LLM condition. This subdued response was frequently attributed by some participants to the relatively more generic nature of the pre-scripted stories, which they felt did not adequately reflect the applicability of the narratives to their own personal experiences.
Participants in the LLM condition felt that their stories were more effective at inspiring solutions. They attributed this success to the stories’ relatability and the fact that they ended on a positive note, offering a sense of hope and direction. As CL37 explained:
Just having a story that reflects mine with a happy ending is something that is important to me and helps provide guidance.
Participants reported learning new techniques through the storytelling process, which helped them reframe their challenges or cope with their emotions more effectively. For instance, CL51 shared how they discovered specific methods for managing anxiety in social situations, particularly when interacting with people with whom they were not comfortable. SL81 echoed the benefits of reading about tangible actions, saying:
The reminder that working towards my goals is something that I can do to increase my self-worth was good. It’s easy to get lost in those negative emotions, and having something tangible to do about it is better than just hearing that I shouldn’t think those things or compare myself to others.
Participants expressed that this reflection process involved more than just passive reading. The LLM-enhanced stories helped participants to reflect on their personal struggles through the lens of the story and to actively engage with the suggested solutions. By presenting coping mechanisms embedded within the context of the narrative, participants felt the stories allowed them to see how similar strategies might apply in their own lives. The stories not only illustrated relatable challenges but also modeled potential solutions in a way that felt accessible and actionable. SL64 commented:
Plainly writing out the advice key message [[by the story]] was helpful since, by relating to the situation, one would feel like they’re being helped by someone in real life as well.
In summary, participants in both groups endorsed the encouragement to actively engage in reflection through writing. However, LLM-enhanced stories were perceived as better at inspiring solutions by depicting progress and providing hope.
4.2.3. Dual Aspects of Relatability
Participants in the Non-LLM condition occasionally reported that their stories were not relatable enough to foster a sense of connection with the story characters. Participants in the LLM condition, on the other hand, noted that their stories fostered a sense of connection with the story characters, as participants recognized that others were experiencing similar challenges. This sense of shared experience provided reassurance, helping participants feel less isolated in their struggles. For example, SL30 expressed:
I like knowing that it’s a shared experience and you’re not alone in this. …it’s reassuring to know that other people struggle with it too.
However, participants like CL3, CL22, and SN66 mentioned that their stories seemed overly tailored to their specific responses, which created a sense of artificiality. They felt that using the exact examples they had provided actually hampered the impact of the stories since it felt improbable that someone else would have experienced the exact same situation. These participants suggested that, in attempting to customize the stories to fit their context, the narratives may have activated suspicion about the story generation process. These observations suggest a nuanced balance is necessary: while insufficient relatability can leave participants feeling disconnected, overly personalized narratives may seem improbable or forced.
4.2.4. Negative Perception of AI Contributing to Reduced Impact
Participants in the LLM condition also suggested that the potential involvement of AI in the story-generation process might activate some negative associations that could undermine their ability to benefit. Participants like CL86, SL23, and SL57 expressed that they were uncomfortable with the idea of AI-based systems providing support for mental wellbeing and expressed a preference for direct human interaction. Participants even pointed out potential risks regardless of the relatability of the AI-generated story. SL57, for example, highlighted this concern:
I don’t think an AI-generated response is going to be helpful in any way. It actually just makes me more depressed thinking about how advice and therapy, a very human task, are being delegated to the 1s and 0s.
Interestingly, some participants in the Non-LLM condition mistakenly believed that AI had been used to generate the stories, and they expressed similar concerns to those in the LLM group. This suggests a broader skepticism among participants about the role of AI in content production.
5. Discussion
Our study demonstrates how LLMs can significantly enhance human-generated content so that it is more relatable and impactful to new audiences. We found that these enhancements improved the utility of the stories as interventions for managing negative thoughts among young adults. These findings add to the growing body of literature on human-LLM collaboration (Chan et al., 2023; Denny et al., 2023; Bhattacharjee et al., 2024b; Guingrich and Graziano, 2023; Sharma et al., 2023a), which seeks to utilize and expand upon existing support mechanisms and content, while shining a light on the evaluation and scrutiny required to support the design of such tools for DMH (Bhattacharjee et al., 2024b; Korngiebel and Mooney, 2021; Demszky et al., 2023).
In our discussion, we begin by addressing how our findings address our research questions. We then discuss the design implications of our results, specifically focusing on future DMH interventions and storytelling applications.
5.1. Key Insights
5.1.1. RQ1: LLM-enhanced Stories in Exhibiting Narrative Qualities
Our research highlights the promising capabilities of LLMs to dynamically generate content that can enhance the quality of narratives in DMH interventions. Similar to prior works (Sharma et al., 2023b; Kim et al., 2024; Bhattacharjee et al., 2024b), the LLM-enhanced stories in our study processed open-ended inputs from participants and adapted the stories to their unique contexts. This capability underscores the potential of LLMs in crafting personalized stories that can support individuals in managing their specific barriers to psychological wellbeing.
Our findings reveal that the LLM-enhanced stories in our study enhanced most of the narrative qualities we investigated. The LLM-enhanced stories were effective in communicating clear and actionable takeaways, which participants noted as essential for translating the stories’ lessons into their own lives (Bhattacharjee et al., 2022; Dyer, 1982). These stories were perceived to excel in promoting reflection, as explicit suggestions within the narratives encouraged readers to contemplate potential changes in their thinking patterns (Bhattacharjee et al., 2024a; Kocielnik et al., 2018). Additionally, the LLM-enhanced stories balanced positivity with realistic portrayals of struggles. The stories often charted a structured progression from challenging life scenarios that resonated with participants’ experiences to outcomes showing small yet meaningful progress (Bhattacharjee et al., 2022).
The LLM was able to render all of these benefits while still being comparable in authenticity to human-written stories, which likely reflects that the LLM effectively mirrored participants’ life experiences and that it drew on human-written example stories. However, while not statistically significant, we also note that the trend in the difference between the two conditions, favored the Non-LLM conition. Our qualitative data also revealed concerns about authenticity in the LLM condition, which could potentially reflect specific word choices made by the LLMs, participants’ familiarity with typical responses made by popular chatbots (Ariyaratne et al., 2023), and the occasional mechanical tone inherent in AI-generated content (Cheng et al., 2022). Recent studies have also examined text generated by ChatGPT across various contexts, noting an increased and repetitive occurrence of certain words such as ‘delve’, ‘intricate’, and ‘underscore’ (Kobak et al., 2024). The frequent use of certain relatively formal words or phrases may contribute to compromising the perceived authenticity of LLM-generated stories.
5.1.2. RQ2: LLM-Enhanced Stories as an Intervention to Help People Manage Negative Thoughts
Our findings suggest that LLM-enhanced stories were relatively more effective in helping individuals reduce their belief in their negative thoughts, aligning with prior literature that underscores the benefits of structured reflection activities for mental health management (O’Leary et al., 2018; Bhattacharjee et al., 2024a, 2022). These stories facilitated reflection on personal thought patterns, allowing participants to draw parallels between their own challenges and those faced by the characters within the narratives (Akram et al., 2020). The LLM-enhanced stories appeared to promote diffuse sociality (Burgess et al., 2019), reducing feelings of isolation by showing that others may also experience similar struggles. Participants noted that this sense of connection reassured them that they were not alone in their challenges.
Narratives that depicted progress, even if minor, provided participants with hope and a sense of direction. Many participants utilized the stories as an opportunity to critically assess and reevaluate their thought patterns. This reflection encouraged them to consider possible changes such as adopting new perspectives on their experiences and exploring alternative actions (Bhattacharjee et al., 2024a), aligning with previous research on the benefits of peer narratives (Bhattacharjee et al., 2022; Akram et al., 2020; Lipsey et al., 2020).
The LLM-enhanced stories also highlight the role of affective empathy in DMH interventions (Sharma et al., 2023a, 2021; Picard, 2000). Affective empathy describes an entity’s ability to perceive, respond to, and convey emotions (Picard, 2000). The LLM-enhanced stories demonstrated this capability by dynamically adapting the content to reflect how users felt about the challenges they encountered. Participants particularly valued how the stories acknowledged the characters’ struggles, which resonated deeply with their own experiences. They reported significant benefits, such as the ability to reevaluate feelings of failure or explore alternative solutions — outcomes that align with the recognized advantages of empathetic interactions in DMH tools (Sharma et al., 2023b; Ardenghi et al., 2024; Pereira and Díaz, 2019).
Finally, relatability was often cited as a key factor enhancing the impact of the LLM-enhanced stories within our study, underscoring the importance of tailoring DMH content to individual contexts (Bhattacharjee et al., 2023b; Baumel et al., 2019). Participants frequently noted that the stories reflected their own experiences, aiding them in visualizing how the advice could be applicable to their own lives (Bhattacharjee et al., 2022). However, excessive relatability detracted from the story’s applicability, as it gave the impression of the story being unrealistically tuned to an individual’s unique situations. These comments bear similarities with the uncanny valley effect (Seyama and Nagayama, 2007) encountered in robotics, which entails discomfort when replicas appear almost, but not quite, human. In the context of DMH, excessively tailored stories can also seem similarly artificial, potentially causing participants to view them as forced and potentially undermining their effectiveness.
5.2. Design Recommendations
Our findings guide us to make several design recommendations. We describe them below.
5.2.1. Balance Between Relatability and Avoiding Implausibility
Our findings highlight a delicate balance in crafting stories that are relatable enough to engage and inspire without crossing into implausibility. Future interventions could address this challenge by creating stories that reflect key aspects of a person’s experience without mirroring their situation too closely, which could otherwise lead to a loss of credibility (Seyama and Nagayama, 2007).
One approach could be to offer users a curated set of story options with varying degrees of relatability. For instance, instead of presenting just one narrative, participants might choose from a range of stories that share similar emotional challenges but differ in specifics like context, background, or tone (Bhattacharjee et al., 2022; Meyerhoff et al., 2024). This would allow users to select a narrative that aligns with their preferred level of resonance. Another approach could involve embedding multiple branching points within stories, where users could navigate different pathways as the narrative unfolds (Bhattacharjee et al., 2022). For example, a user could choose whether the protagonist tackles their challenges by seeking external support or by relying on personal coping strategies. This interactive experience might offer users control over how much the story connects with their own experiences, making the process feel more dynamic and tailored.
5.2.2. Refining Tone and Wording of AI-generated Content
While LLM-enhanced stories demonstrated promising potential as narrative interventions, some participants expressed concerns about the writing style and tone, which they felt diminished the overall impact. This issue arose from the use of certain phrasing or the participants’ familiarity with the way ChatGPT responds, making the stories feel less authentic. The tone of the responses was frequently mentioned, with some participants finding the tone to be too mechanical lacking the personal touch necessary for conversations around managing negative thoughts (Cheng et al., 2022). Others noted that the overly optimistic tone in certain stories did not always align with or acknowledge the severity of their struggles, which, in turn, made them feel that their challenges were being minimized (Haugeland et al., 2022).
To address these issues, future DMH interventions could focus on refining the tone and wording of LLM-generated stories. For example, incorporating more variability in sentence structures and word choice could make the narratives feel less formulaic and more natural (Park et al., 2023). This enhancement could be achieved by explicitly instructing LLMs through customized prompting, experimenting with the temperature parameter to introduce variability, or fine-tuning LLMs using stories that contain diverse structures. Another approach could be to vary the emotional tone of stories depending on the intensity of the challenges being addressed, ensuring the tone is empathetic without being overly upbeat (Haugeland et al., 2022). Additionally, allowing for more ambiguity or open-ended conclusions in some stories might prevent them from feeling overly prescriptive, giving participants room to reflect on their own situation rather than feeling as though they are being given an unrealistic solution (Bhattacharjee et al., 2022). However, all of these refinements should be carefully tested to tailor the tone, structure, and ambiguity to different user needs.
5.2.3. Structured Reflection Activities
Participants in our study expressed appreciation for the reflection activity embedded within the stories, noting that such exercises helped them reevaluate their thoughts and emotions and potentially identify solutions to their problems. However, the reflection activity in our study was limited to a single question. Future studies might consider drawing inspiration from guided reflection activities, which involve a sequence of targeted questions focusing on specific aspects of one’s thoughts (O’Leary et al., 2018). These multi-question formats allow for a more structured presentation of thoughts, capturing nuances that a single question might miss. Incorporating such detailed questioning can enhance the benefits of LLM-enhanced stories by providing deeper insights into their thoughts and challenges, and potentially helping find ways of solving them. (Bhattacharjee et al., 2024a; Sharma et al., 2023b). However, it is important to note that answering these more detailed questions requires more time and effort from users. Therefore, careful consideration should be given to the timing and frequency of these questions to maintain a balance between depth of engagement and participant burden.
5.2.4. Improving Peer Support and Crowd-Powered Platforms
In the past, peer support and crowd-powered platforms have shown promising potential in providing wellbeing support, yet they face several inherent challenges that could be alleviated by an approach like ours. For example, the act of sharing and writing content can be daunting; LLMs can assist crowdworkers and peers by providing example stories and offering writing support to edit and restructure users’ own experiences (Kim et al., 2024; Reza et al., 2023). Additionally, some individuals may be reluctant to open up to real people. In these cases, the ability of LLMs to craft highly personalized narratives in a completely private manner — without the need to share personal details — could significantly increase their appeal, particularly for more inhibited users. Nonetheless, it is crucial to position LLMs as supportive tools that enhance the human-driven process and maintain the essential human element in peer support intact.
5.2.5. Considerations for People’s Attitude Toward AI
While our study did not specifically measure attitudes towards AI, some participants’ comments suggest that such attitudes may influence their perceptions of LLM-enhanced stories. Some of our participants expressed resistance to the idea of AI-based tools providing support for managing negative thoughts while others did not, suggesting a wide range of expectations regarding AI’s role in personal interactions (kaya2024roles; Liu et al., 2022). This variation suggests that future research could benefit from directly measuring attitudes toward AI (Grassini, 2023) and examining how these attitudes shape preferences for LLM-generated stories. Additionally, future DMH tools could incorporate a feature allowing users to express their attitudes toward AI during registration. This mechanism could enable the system to adjust its responses to better match individual expectations and comfort levels with AI technology. Moreover, AI systems could dedicate time to build trust gradually, rather than immediately offering wellbeing support, which some individuals might find uncomfortable (O’Dell et al., 2022).
5.3. Limitations and Future Works
Our study has several limitations that merit attention. First, the engagement of participants with the stories was limited to a one-time survey activity. While brief reflection activities like these are often tested in singular instances, the long-term effects of such interventions are also critical (Bhattacharjee et al., 2024a). A longitudinal deployment could uncover factors such as time, user mood, emotion, and social context that might influence engagement with the stories (Bhattacharjee et al., 2023b). Future research should explore these dimensions.
Secondly, our study population was confined to young adults. While some findings from this demographic might be generalizable to other groups, conducting similar studies with different populations could reveal new insights into their preferences and responses. Future studies should also consider including individuals with clinically elevated levels of depression and anxiety symptoms, keeping in mind the heightened need for risk management processes and oversight in such research.
Lastly, our study relied primarily on user perceptions of the LLM-enhanced stories — their views on various qualities of the stories and their effectiveness as interventions. That is, we did not evaluate objective qualities of the content itself. While we constrained the LLMs to work within the framework of previously established narratives supporting evidence-based psychological strategies, future research should assess adherence to these strategies and other inappropriate content that may appear in generated stories, potentially involving clinicians in the evaluation process to ensure the accuracy and appropriateness of the content.
6. Conclusion
LLMs offer promising capabilities to dynamically adapt and generate stories that can resonate with the unique challenges individuals face. Motivated by this potential, our study leveraged LLM to enhance existing human-written stories from past DMH interventions, comparing the LLM-enhanced versions with their original counterparts. We conducted an experiment with young adults, including crowdworkers and students, and observed that LLM-enhanced stories were perceived to do better than human-written ones in conveying key takeaways, promoting reflection, and reducing belief in negative thoughts. These narratives were not only perceived as more relatable but also maintained a level of authenticity comparable to human-written stories. Our findings illuminate potential areas for future work on LLM-enhanced narrative interventions, such as balancing relatability with avoiding implausibility and refining the AI’s tone and writing style. We believe our work marks a significant step towards enhancing DMH interventions with emerging AI technology, enabling them to dynamically adapt based on user challenges.
Acknowledgements.
This work was supported by grants from the National Institute of Mental Health (K01MH125172, R34MH124960), the Office of Naval Research (N00014-18-1-2755, N00014-21-1-2576), the Natural Sciences and Engineering Research Council of Canada (RGPIN-2019-06968), and the National Science Foundation (2209819). In addition, we acknowledge a gift from the Microsoft AI for Accessibility program to the Center for Behavioral Intervention Technologies that, in part, supported this work (http://aka.ms/ai4a). Ananya Bhattacharjee would also like to acknowledge support from the Inlight Research Fellowship, the Schwartz Reisman Institute Graduate Fellowship, and the Wolfond Scholarship in Wireless Information Technology.
Appendix A Examples of Stories
Table 5. Non-LLM stories accompanied by LLM-enhanced versions
User’s Negative Thought User’s Provided Description Non-LLM Story LLM Story
I always feel really down on myself I feel disappointed in myself. There are many goals I set that I have not reached such as getting on the Dean’s List scholar and making new friends in my first year. University has been a really tough transition for me since I moved abroad and I feel super lonely most of the time. I want to focus on my grades, but then I won’t have enough time to make friends in class or through extracurricular activities. I am also quite introverted so it’s hard for me to talk to new people. I envisioned a different university life. I just often feel down that none of my expectations are met. I am really really hard on myself sometimes. I just keep fixating on and thinking about things I have done wrong.
But I am starting to be a bit more forgiving with myself because this is not how I want to feel and not how I want to treat myself. One of the ways I’ve been doing this is by telling myself a few positive affirmations every day. I’ll tell myself over and over again things like “it’s okay to make mistakes,” “it’s okay not to feel happy,” “it’s my right to be able to tell others no,” or “I am trying my best.”
I know it seems super silly and can feel a little awkward at times, but I’ve noticed that sometimes I’ll say one of the affirmations in response to some of my bad thoughts about myself, so it’s become more like a habit. They’re just simple, true things that I sometimes lose sight of.
Being deliberate about the things we tell ourselves can make a difference in how we feel.
Please answer this reflective question: What simple, true things do you sometimes lose sight of?
Life hit hard when I moved to university, in a city miles away from home. Just like you, I was quite introverted and struggled with balancing academics and social life, multiply that with the pressure of getting on the Dean’s list scholar. Loneliness became my consistent companion.
Then a friend told me about the power of positive affirmations. I dismissed it at first, but as I grew lonelier, I thought, why not give it a try? So, I started practicing. I would tell myself things like ’it’s okay to make mistakes,’ ’it’s okay to feel lonely’ or ’I’m not failing, I’m learning.’ It felt awkward and silly in the beginning, but slowly, I noticed a shift in how I felt about myself.
Every time I started to feel down on myself, a self-affirmation would pop into my mind. I realized it was not about being perfect, but about focusing on my growth, my journey. This simple act of self-love still helps me during my bad days at university, and it’s worth trying.
Please answer this reflective question: What kind phrases can you say to yourself during moments of self-disapproval or disappointment?
A lot of things could go wrong in the future While things are fine in my life right now, I can’t stop thinking about how a small incident could make me lose everything. I could make a mistake at work that gets me fired, and lose the ability to provide for my family. How will I feed myself or my daughter and make rent? Currently, I am in the process of applying to grad school and I am worried about not getting in. I feel that if I don’t get in, I will be very disappointed in myself. I find myself frequently going through various scenarios about all the things that could go wrong.
I know this is only hurting me, so I am trying to be more self-compassionate. I am trying to be as nice to myself as I would be to a friend who was applying to grad school.
If I were talking to a friend, I’d probably say something like, “You are smart and successful already. I hope it works out, but if it doesn’t you can reapply or find another path that will work for you.” It’s something I’m trying to remind myself so I can relax a little.
Self-compassion is all about giving the same kindness to ourselves we give to others.
Please answer this reflective question: How can you be kind to yourself next time you’re caught up in worry?
As I lay awake one night consumed by a whirlpool of thoughts about the potential problems I might face at work the next day, I suddenly realized how much I’d been tormenting myself with the ’what ifs’. I was so enveloped in the fear that a minor mistake could unravel everything, it took over my ability to live in the present.
That’s when I decided to start treating myself with more compassion, as though I were a dear friend stuck in the same situation. I imagined what I’d say to them, ’You are competent and have tackled obstacles before, surely you can do it again. Even if the unforeseen happens, you’ll find a way to pull through it.’ I began repeating this to myself, becoming my own advocate.
Over time, I learned to cope better with my fears. Instead of showcasing my life’s fragility, these anxieties became reminders that even if things go wrong, I have the capacity to bounce back.
Remember, self-compassion is giving the same kindness to ourselves that we extend to others. It’s not an overnight transition, but eventually, it eases the burden of constant worry.
Please answer this reflective question: How can you apply self-compassion in your daily life when facing the fear of uncertain future outcomes?
I am struggling more than I should I’ve been struggling financially, and it feels like I should have it figured out by now. I’m turning 25 this year, and it’s hard to shake the feeling that I’m too old to still be barely making enough to cover rent. I wish I could be on top of my finances, instead of constantly worrying about money and just scraping by. I have always struggled in life: struggled to fit in, struggled to sleep, struggled to read. It just seemed like in all the ways that people are successful, I struggled. I really started to feel down about my struggles in high school but it wasn’t until my undergraduate years that I started seeing a professional.
I was lucky to be in school and to have counseling options available to me. I know there are some stigmas attached to seeing a professional, but I just needed someone unbiased, professional, and who could perhaps really help me understand myself better. The best thing that came out of that experience is that I continued to see professionals after college, whenever I felt stuck or was struggling, or just needed someone to listen. I don’t see a therapist every week anymore, but I know I can find help if I need it, and it’s a relief.
Sometimes, we need to reach out to others to get a handle on things. Whether that’s a professional or someone else.
As a bright-eyed 24-year-old, life felt like a jigsaw puzzle I was yet to solve, and finances, ironically the key corner piece. As days turned into weeks, and weeks into months, my wallet seemed to grow lighter, and the weight of my worries heavier. I was on the brink of turning 25 and the thought of barely making ends meet felt heavier than ever.
One day, out of the blue, I remembered how eager I was in college to learn and grow from my peers and professors. It was all about sharing and collaboration then. Why had I stopped reaching out? Was it fear? Was it the stigma? These questions pushed me beyond my comfort zone and I found myself venturing into the world of professional help.
I won’t purport that overnight my troubles were resolved. But seeking help acted as a beacon of hope, gradually illuminating the darker corners of my finance woes. Years later, I still reach out to professionals when I find myself in a rut - not with the expectation of instant resolutions, but to gain perspective and feel understood.
Please answer this reflective question: Can you think of a time when reaching out to someone helped you break out of a rut? Relying on others may sometimes make you feel vulnerable, but sometimes it is in this vulnerability that we find our greatest strengths. There’s no shame in seeking help, after all, it’s okay to not have all the pieces together. We’re all a work in progress.
Please answer this reflective question: Does this narrative prompt you to reach out for professional help, or remind you of a time when doing so was the best decision you took?
I haven’t achieved as much as other people have I feel as though I haven’t accomplished much compared to other 23 year olds, and it sometimes makes me feel horrible and depressed when I compare myself to those who live different lifestyles. Ever since graduating high school 6 years ago and seeing everyone that I grew up with accomplish their goals and start to have their own families, I compare where my life is at in regards to where they’re at. I tend to be really hard on myself and never feel like I’m doing well enough and am good enough.
When the pandemic started, I started working from home, and with places being closed, I was home 24/7. With no way of really ignoring my feelings, I hit an extremely low point. I made a Facebook post that was essentially a cry for help because I just didn’t know what to do. Ultimately, I’m glad that I did.
Six years post-high school and I kept grading myself by my friends’ life charts, their flourishing careers, their growing families. I had painted a grueling picture of my life, always falling short of what they had achieved.
When the world quietened due to the pandemic, and I was confined at home, these comparisons became deafening echoes in my solitude. One harsh day, I succumbed to it and shared an undeclared plea for help on Facebook, a confession of my perceived underachievement. I was at the edge of despair.
Then occurred a shift I hadn’t anticipated.
A very dear friend of mine reached our privately to express he cares about me. I have been talking to my friend more often, who shared with me that he has gone through similar things as well. It’s helped me see that acknowledging how I feel can help.
Our conversations have been a soothing balm, a beacon of shared experiences, providing comfort and perspective. So, remember, you’re not alone in this tumultuously difficult journey. We all have stories, often unspoken, all challenging and unique.
If we only see the good sides of people’s lives that they share, it’s easy to forget how much other people struggle too.
Please answer this reflective question: What do you wish people would be more open about?
A dear friend reached out with kind words, expressing his concern and shared his battles with similar circumstances. His candidness evoked an undeniable truth - our social media lives are mostly highlight reels, the failures and struggles often invisible. We, unknowingly, harshly measure our entire lives to others’ well-curated feeds.
Please answer this reflective question: Have you reached out to a friend when you felt overwhelmed with self-doubt and how did that conversation make you feel about your own journey?
The people I have counted on have let me down I feel really let down by my partner. I thought I could count on them, but when I needed them most, they weren’t there. It’s hard to shake the feeling of betrayal, like I don’t matter as much to them as they do to me. I had to become independent at a very young age as I realized there was no one I could depend on. This has made it hard for my partners in relationships, as I don’t know how to really let people in. I want to, but I’ve seen what happens when you let someone in. It can be devastating. The first time I completely trusted and depended on someone, I was left shattered for 7 years.
I was thrust into independence at an early age, having realized that relying on others often led to disappointment. This developed a hardened shell around me, especially in romantic relationships where the fear of vulnerability reigned. I yearned for connection, yet the thought of opening up was paralyzing. Once I did open up, the outcome was a heartbreak that took me seven years to recover from.
Over that time, I suffered from chronic depression. I pushed people away in an attempt to protect myself.
I have come to accept that we can’t hate people for not sticking with us when they want to do something else. Negative experiences are a part of life. We can either let them swallow us, or we can push forward. I still cry sometimes, and some situations get the best of me.
There will be hard times, but you’ve gotten through them before and will again.
Please answer this reflective question: What helps you get through it when people let you down?
During those years, a shroud of chronic depression enveloped me. I pushed everyone away, building an even thicker protective barrier around myself.
Over time, acceptance found its way into my thought process.
I acknowledged that life is a blend of positives and negatives, and people’s decisions are governed by their own desires. We can’t bind them by our expectations. Yes, there are moments of weakness when I still surrender to tears.
But with every sunrise, there’s an opportunity to pick myself up and march on. Sometimes, surviving the whole day seems daunting, but taking it one-minute at a time is manageable. Through all life’s trials and tribulations, remember this: you’ve faced hard times before and emerged victorious.
Please answer this reflective question: What strategies do you find useful to cope when you feel let down by those you count on?
I feel too shy or anxious when I’m in social situations I feel overwhelmed and awkward in social situations, like I don’t know how to fit in or what to say. The anxiety just takes over, and I end up feeling invisible or out of place. I am a shy person. It’s part of who I am. I sometimes feel bad about this, like I have lost out on experiences because of being shy and not showing my true self to the world. I’ve often found myself tangled in the webs of social anxiety, like a silent character adrift in a sea of voices. There’s this persistent whisper in my psyche, convincing me that everyone in the room is scrutinizing each slip, each stumble, and each stutter.
It’s exhausting, and I wish I could relax and be myself around others. While I know I am as worthy of having a voice as anyone else, there’s this worry in my head that others will judge me.
What has helped me is to argue against my worries when I have them. I try to push back when the voice in my head says “they’re judging you”.
Plus, I have a lot of positive qualities. When people get to know me, they usually like me. My worries are not based on the reality of how people think. Remembering this has helped me feel a little more confident when I’m meeting someone new.
It’s easy to forget that other people are often pretty focused on themselves, and don’t spend as much time noticing and judging us as we might think.
Please answer this reflective question: What strategies work for you when you feel shy or socially anxious?
It’s a burden, and all I yearn for is to let go, to be the authentic “me” amidst the crowd.
I found solace in challenging this distressing voice in my head. Every time it chanted “they’re evaluating you”, I would confront it, argue against it.
I started to realize that people seldom ponder about my awkwardness as much as I imagine. Let’s face it, we each have our own worlds to worry about.
Interestingly, I’ve come to embrace my introverted self as part of my identity. I’ve learned that this part of me possesses its own unique charm, and as people venture closer, they often appreciate my depth and catalogue of thought. This understanding has fostered a spark of confidence when meeting new faces.
In the vast universe of their own thoughts and worries, people rarely fixate on our quirks or mishaps. Embracing this fact has brought me a step closer to accepting my social struggles and actively working on overcoming them.
Please answer this reflective question: How do you combat your worries when you find yourself feeling socially anxious or shy? Are there certain affirmations or thought patterns that seem to help you most?